OpenStudy (anonymous):

Together, Tim and Alan can unload a truck 8 in minutes. When Tim works alone, it takes him 12 minutes longer than Alan to unload the truck. How long does it take each person to unload the truck alone?

OpenStudy (anonymous):

@superhelp101

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

i got an answer by trial and error

OpenStudy (anonymous):

maybe we could do some algebra to get it , but i guessed and checked you want some algebra or my guessing?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

tim 18?

OpenStudy (anonymous):

sure which one? i got \(12\) and \(24\)because \[\frac{12\times 24}{12+24}=8\]

OpenStudy (anonymous):

so whose 24

OpenStudy (anonymous):

tim? and alan 12

OpenStudy (anonymous):

i don't know the slower one has rate 12

OpenStudy (anonymous):

When Tim works alone, it takes him 12 minutes longer than Alan to unload the truck. Tim is slower his rate is 24, alan's rate is 12

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

that is my answer i would be $14 on it but not more

OpenStudy (anonymous):

your right!

OpenStudy (anonymous):

thanks!:)

OpenStudy (anonymous):

imagine trial and error wins again

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

lol!!